# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v5

@@ 2.1 Problem Set-Up
where $0\div 0$ is defined to be 0. Standard design methods also control power with respect to some data generation process $\Psi$. The goal of our proposed method is to recommend a suitable $(n,\boldsymbol{\gamma})$ combination such that the criteria for the Bayesian FDR in (1) and average power in (3) are approximately attained. We formally describe how our method attains these criteria as the recommended value of $n\rightarrow\infty$ in Section 4. In the next subsection, we explain why standard methods to find suitable $(n,\boldsymbol{\gamma})$ combinations are very computationally intensive and how we improve upon standard design procedures.
@@ 4 Economical Assessment of the FDR and Power
However, we recommend defining $\Psi$ as a mixture of various submodels, each of which gives rise to a different $\boldsymbol{\theta}_{r}$ value, to accommodate uncertainty about which hypotheses are false. This submodel approach is demonstrated in Section 5. In that case, the process in Lines 10 to 16 can be modified. We instead split the logits of the posterior probabilities for each $n$ value into subgroups based on their submodels in $\Psi$ before constructing the linear approximations. We use this set of linear approximations to obtain the optimal sample size $n_{2}$ and its associated decision thresholds in Line 17. Corollary 1 details how the Bayesian FDR and average power for the design defined using the $(n,\boldsymbol{\gamma})$ combination returned by Algorithm 2 respectively converge to $q$ and $1-\beta$ as the recommended value of $n\rightarrow\infty$.
@@
Let the conditions for Theorem 1 be satisfied. Let the Bayesian FDR and average power be defined as in (1) and (3), respectively. Consider a design with the $(n,\boldsymbol{\gamma})$ combination recommended by Algorithm 2. As the recommended $n\rightarrow\infty$, this design is such that
$\left\lvert\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\frac{V}{V+S}\right]\right\}-q\right\rvert\xrightarrow{P}0~~\text{and}~~\left\lvert\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\frac{S}{T}\right]\right\}-(1-\beta)\right\rvert\xrightarrow{P}0.$
We prove Corollary 1 in Appendix C of the online supplement. While Algorithm 2 leverages asymptotic results, its linear approximations are constructed by simulating sampling distributions of posterior probabilities at finite sample sizes $n$. With finite sample sizes, Algorithm 2 thus recommends $(n,\boldsymbol{\gamma})$ combinations such that the Bayesian FDR approximates $q$ and average power approximates $1-\beta$. In Section 5, we consider the performance of this algorithm for an example based on a real A/B test involving 5 metrics.

### v6

@@ 2.1 Problem Set-Up
where $0\div 0$ is defined to be 0. Standard design methods also control power with respect to some data generation process $\Psi$. The goal of our proposed method is to recommend a suitable $(n,\boldsymbol{\gamma})$ combination such that the criteria for the Bayesian FDR in (1) and average power in (3) are approximately attained. We formally describe how our method approximately attains these criteria in Section 4. In the next subsection, we explain why standard methods to find suitable $(n,\boldsymbol{\gamma})$ combinations are very computationally intensive and how we improve upon standard design procedures.
@@ 4 Economical Assessment of the FDR and Power
However, we recommend defining $\Psi$ as a mixture of various submodels, each of which gives rise to a different $\boldsymbol{\theta}_{r}$ value, to accommodate uncertainty about which hypotheses are false. This submodel approach is demonstrated in Section 5. In that case, the process in Lines 10 to 16 can be modified. We instead split the logits of the posterior probabilities for each $n$ value into subgroups based on their submodels in $\Psi$ before constructing the linear approximations. We use this set of linear approximations to obtain the optimal sample size $n_{2}$ and its associated decision thresholds in Line 17.
We provide a theoretical guarantee for the performance of Algorithm 2 below. Note the recommended sample size $n$ is a function of the bound $q$ for the Bayesian FDR, the target value $1-\beta$ for average power, the probability model $\Psi$, and the prior $p(\boldsymbol{\eta})$. To allow the recommended $n$ value to approach $\infty$ in our theoretical results, we introduce a sequence for average power $(1-\beta_{n},\dots)$ that increases to 1. Corollary 1 details how the Bayesian FDR and average power for the design defined using the $(n,\boldsymbol{\gamma})$ combination returned by Algorithm 2 with inputs $q$ and $1-\beta_{n}$ respectively converge to $q$ and $1-\beta_{\color[rgb]{.75,0,.25}n}$ as the recommended value of $n\rightarrow\infty$.
@@
Let the conditions for Theorem 1 be satisfied. Let the Bayesian FDR and average power be defined as in (1) and (3), respectively. Consider a design with the $(n,\boldsymbol{\gamma})$ combination recommended by Algorithm 2 with inputs $q$ and $1-\beta_{n}$. As the recommended $n\rightarrow\infty$, this design is such that
$\left\lvert\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\frac{V}{V+S}\right]\right\}-q\right\rvert\xrightarrow{P}0~~\text{and}~~\left\lvert\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\frac{S}{T}\right]\right\}-(1-\beta_{\color[rgb]{.75,0,.25}n})\right\rvert\xrightarrow{P}0.$
We prove Corollary 1 in Appendix C of the online supplement. While Algorithm 2 leverages asymptotic results, its linear approximations are constructed by simulating sampling distributions of posterior probabilities at finite sample sizes $n$. With finite sample sizes that are sufficiently large, Algorithm 2 with fixed inputs $q$ and $1-\beta$ thus recommends $(n,\boldsymbol{\gamma})$ combinations such that the Bayesian FDR approximates $q$ and average power approximates $1-\beta$. In Section 5, we consider the performance of this algorithm for an example based on a real A/B test involving 5 metrics.